OpenStudy (anonymous):

solve: log(x+4)=logx + log4 show work

OpenStudy (asnaseer):

do you know the log rules for adding two logs?

OpenStudy (asnaseer):

if not, have a look here first: http://www.mathsisfun.com/algebra/exponents-logarithms.html and then try to attempt your question. If you still have trouble then post where you are getting stuck here and I'll try and help you out.

OpenStudy (anonymous):

yes i know the rules. i got to: (x+4/4x)=0 ...is that the right track?

OpenStudy (asnaseer):

you went one step too far... you probably got to this step first: log(x+4) = log(4x) - correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then subtracted to set it equal to zero

OpenStudy (asnaseer):

ok, so you have log(something) = log(something else) which implies something = something else

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

but i need an actual answer.... it's supposed to be 12

OpenStudy (asnaseer):

you got to this: log(x+4) = log(4x) which implies: x + 4 = 4x now just solve to find x

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

yw :)